Solve the following quadratic programming problem(Wolf’s method)

Maximize         Z = 2x1+x2-x12,

Subject to        2x1+3x2 ≤ 6,

                          2x1+x2 ≤ 4,

               x1, x2≥ 0


1
Expert's answer
2021-11-19T01:24:48-0500

Step-1 :

The problem is converted to canonical form by adding slack, surplus and artificial variables as appropiate.


"2x_1+3x_2+s_1=6\\2x_1+x_2+s_2=4"



The problem is converted to canonical form by adding slack, surplus and artificial variables as appropiate.

"Z'-2x_1-2x_2+x_3=0"

"2x_1+3x_2+s_1=6\\2x_1+x_2+s_2=4"



Now represent the new system of constraint equations in the matrix form

"\begin{bmatrix}
 1 & -2&-1&1&0&0 \\
 0 & 2&3&0&1&0\\0 & 2&1&0&0&1
\end{bmatrix}\begin{bmatrix}
 Z'\\
 x_1\\x_2\\x_3\\s_1\\s_2
\end{bmatrix}=\begin{bmatrix}
 0 \\
 6\\4
\end{bmatrix}"

"or\\" "\begin{bmatrix}
 -1 & -c \\
 0 & A
\end{bmatrix}\begin{bmatrix}
 z \\
 x 
\end{bmatrix}=\begin{bmatrix}
 0\\
 b
\end{bmatrix}\geq0"


"where e=\u03b2_0,a_3=\u03b2_1,a_4=\u03b2_2"


"Step\ 2:The\ basis \ matrix\ B_1\ of\ order (2+1)=3 can\ be\ expressed\ as\\ \beta_1=\begin{bmatrix}
 \beta_0 & \beta_1&\beta_2 \\
 
\end{bmatrix}=\begin{bmatrix}
 1 & 0&0\\
 0& 1&0\\0 & 0&1
\end{bmatrix}"





Iteration=1 : Repeat steps 3 to 5 to get new solution

Step-3: To select the vector corresponding to a non-basic variable to enter into the basis, we compute

"c_k-z_k=Max\lbrace{(c_j-z_j)>0;j=1,2}\rbrace"


"=Max\lbrace{-(1^{st}\ row\ of\ \beta_1^{-1})(Columns\ a_j\ not\ in\ basis)}\rbrace"


"=max\lbrace-\begin{bmatrix}
 1 & 0&0
\end{bmatrix}\begin{bmatrix}
 2 & -1& 1 \\
 2 & 3& 0\\2 & 1& 0
\end{bmatrix}\rbrace"


"=Max\lbrace-\begin{bmatrix}
 -2 & -1 &1
\end{bmatrix}\rbrace\\ =Max\lbrace\begin{bmatrix}
 2 & 1 &-1
\end{bmatrix}\rbrace\\ =2 (correspnds\ to\ c_1-z_1)"

Thus, vector "x_1" is selected to enter into the basis, for k=1


Step-4: To select a basic variable to leave the basis, we compute "y_k" for k=1, as follows


"y_1=\beta_1^{-1}a_1=\begin{bmatrix}
 1 & 0&0\\
 0& 1&0\\0 & 0&1
\end{bmatrix}\begin{bmatrix}
 -2\\
 2\\2
\end{bmatrix}=\begin{bmatrix}
 -2\\
 2\\2
\end{bmatrix}and\ X_B=\begin{bmatrix}
 0\\
 6\\4
\end{bmatrix}"


Now, calculate the minimum ratio to select the basic variable to leave the basis


"\frac{x_{\beta r}}{y_{rk}}=min\lbrace\frac{x_{\beta i}}{y_{ik}},y_{ik}\gt0\rbrace\\"

"=min\lbrace \frac{6}{2},\frac{4}{2}\rbrace\\=min\lbrace3,2\rbrace"


"=2(corresponds\ to \frac{x_{\beta 2}}{y_{21}} )"


Thus, vector "S_2" is selected to leave the basis, for r=2


The table with new entries in column "Y_1" and the minimum ratio






The table solution is now updated by replacing variable "S_2" with the variable "X_1" into the basis.


For this we apply the following row operations in the same way as in the simplex method



The improved solution is




Iteration=2 : Repeat steps 3 to 5 to get new solution

Step-3: To select the vector corresponding to a non-basic variable to enter into the basis, we compute


"c_k-z_k=Max\lbrace{(c_j-z_j)>0;j=1,2}\rbrace"


"=Max\lbrace{-(1^{st}\ row\ of\ \beta_1^{-1})(Columns\ a_j\ not\ in\ basis)}\rbrace"


"=max\lbrace-\begin{bmatrix}
 1 & 0&1
\end{bmatrix}\begin{bmatrix}
 0 & -1& 1 \\
 0 & 3& 0\\1 & 1& 0
\end{bmatrix}\rbrace"


"=Max\lbrace-\begin{bmatrix}
 -1 & 0 &1
\end{bmatrix}\rbrace\\ =Max\lbrace\begin{bmatrix}
 -1 & 0 &-1
\end{bmatrix}\rbrace"


"Since\ all\ Z_j-C


_j≥0"


"Hence, optimal\ solution\ is\ arrived\ with\ value\ of\ variables\ as :\\
x_1=2,x_2=0,x_1^{2}=0"


"Max Z=4"